If f(x) = 2|x-1| and g(x) = 3x-10 , then f o g (1) = 12 .


True or false? Give reason.

1
Expert's answer
2021-11-07T16:59:56-0500

As g(1)= 3x-1= 3(1)-1= 2


As fog(1) = f(g1) 

As calculated it would become:


fog(1)= 2|x-1| = 2|2-1| = 2.


Hence, fog(1) = 2